If the time between calls in a call center follows an Exponential Distribution with mean 5 minutes, what is the probability that there will not be another call in the next 10 minutes?


1
Expert's answer
2021-05-04T13:03:54-0400

For the exponential distribution, the mean is

"5 = E[X] = 1/\lambda \; \Longrightarrow\; \lambda = 1/5"


'will not be another call in the next 10 minutes' = 'the next call will be after 10 min'

"P(X>10) =1 - P(X \leq 10) = 1-(1 -e ^{-10 \lambda}) = 1-(1 -e ^{-10/5})= 1-(1 -e ^{-2})= 1 -1 + 0.1353 = 0.1353"

Answer: 0.1353